The vapour pressure of methyl alcohol is 100mmHg at 21.2 degree Celsius. What is it's vapour pressure at 25.0 degree Celsius? (∆Hvap = 38.0kj/mol)
1
Expert's answer
2020-08-26T07:48:46-0400

We solve according to the Clapeyron-Clausius equation

"ln(\frac{P1}{P2})=\frac{\Delta Hvap}{R}(\frac{1}{T2}-\frac{1}{T1})"

P1=100mmHg=0.131579 atm

R=8.314 J/(K×mol)

T2=298.15 K

T1=294.35 K


"ln(\frac{0.131579}{P2})=\frac{38 000}{8.314}(\frac{1}{298.15}-\frac{1}{294.35})"


"ln(\frac{0.131579}{P2})=-0.197905708"


"\frac{0.131579}{P2}=e^{-0.197905708}"

"\frac{0.131579}{P2}=0.820447211"


P2=0.16037473 atm or


1 the physical atmosphere is equal to 760.002211 millimeters of mercury


0.16037473 atm=121.885149 mmHg